# Reasoning Beyond Majority Vote: An Explainable SpeechLM Framework for Speech Emotion Recognition

### v1

@@ 1 Introduction
These limitations are most acute at evaluation time. Majority-only metrics penalize predictions that agree with well-supported but non-majority annotations, thereby underestimating performance on inherently ambiguous utterances and masking where models align with valid minority judgments [5]. To address this, SER research has increasingly embraced soft-label formulations, where the full annotator distribution or uncertainty-aware targets are preserved in training and evaluation [6, 7, 8]. Such methods not only yield higher accuracy on ambiguous cases but also align more faithfully with human variability. Beyond scores, a model requires understanding how systems map linguistic content and acoustic cues (prosody, pitch, energy, timing) to emotions so we can target weaknesses (e.g., lexical shortcuts vs. prosodic evidence). While recent SpeechLLM studies have begun entering the SER domain [9, 10, 11, 12], and some report headline accuracies in technical reports [13], the majority still treat these models primarily as classifiers. This paradigm under-utilizes their generative capabilities, leaving untapped the potential to combine reasoning, explanation, and label uncertainty into a more holistic framework for emotion recognition.
In this work, we introduce an explainable SpeechLM framework that outputs a natural-language rationale alongside the emotion label. A reasoning-capable teacher LLM synthesizes concise explanations grounded in both lexical content and acoustic cues simultaneously. These rationales serve as intermediate supervision signals for a SpeechLM backbone. Given an utterance, the model first emits an ASR transcript to preserve transcription skill, then an emotion decision followed by a brief evidence-based rationale, all within a single coherent response. This structure encourages the model to provide both the emotion prediction and its rationale, improving interpretability while leaving majority-label training unchanged. To our knowledge, this is the first SER study to pair qualitative/quantitative rationale analysis with annotator-aware evaluation on a large in-the-wild corpus.
@@ 2.1 Reasoning Generation
Recent advances in reasoning-capable LLMs have made strong open models widely available. To bridge the gap between speech inputs and categorical emotion outputs, we leverage these models to generate human-readable rationales for each input–output pair in speech emotion recognition (SER). The prompt template used for rationale synthesis is provided in the Prompt Design box, where the \$CONTENT and \$EMOTION mean the transcript and ground-truth emotion labels of the training utterances respectively.
@@
Concretely, we adopt Qwen2.5-Omni as the reasoning agent due to its strong multimodal understanding and explicit reasoning capabilities. For each utterance, we provide the audio signal (encoded into embeddings by audio encoder) concatenated with prompt token embeddings consisting of its transcript and the reference emotion label to generate rationales for training data. Note that unlike prior work, we do not discretize acoustic cues into fixed levels. This strategy enables diverse natural-language descriptions and reflects real-world settings where the target distribution and class granularity are unknown. All rationales are generated exclusively on the training split, ensuring no label information leaks into validation or test evaluation.
@@ 2.2 Supervised Fine-Tuning (SFT)
Table 1: Baseline model comparison on the MSP-PODCAST testing set, and the Macro-F1 score is presented in the table.
@@
11 Q2A-lora-SFT w/ reasoning 36.14% 66.78% 40.15% 59.32% 30.59% 62.55% 33.64% 54.13%
@@ 3.2 Baseline Comparisons
In contrast, our proposed SFT-based model (Model 11) substantially outperforms all zero-shot baselines (including CoT-based models), reaching 36.1% Macro-F1 on MV-labels (open-form) and 30.6% in the closed-form setting. Importantly, these gains are not merely a byproduct of fine-tuning: traditional SFT models trained directly for SER (Models 1–3) achieve only 23–30% on closed-form metrics, highlighting that simply applying supervised training does not guarantee strong performance. Instead, the improvement of Model 10 stems from incorporating rationale supervision during fine-tuning, which enhances both predictive accuracy and interpretability.
3.3 Ablation on Rationales
@@ 3.4 Subjective Test
To assess the reasoning plausibility, we selected $32$ utterances where the model’s predicted label disagreed with the reference label. For each utterance, we formed two candidates including Model-predicted and Gold-referenced (rationale generated by our prompting procedure using the reference label as guidance). Raters marked the preference one, and we report the Gold-vs-Predicted preference rate.
All results are summarized in Figure 1. In Test I, the overall average acceptance rate reached 69.1% (95% confidence intervals(CIs): [58.9%, 79.2%]), with 72.5% for the match condition and 65.6% for the non-match condition. These results indicate that the reasoning generated by our model is perceived as plausible by human raters, even when the predicted emotion does not align with the ground truth.
In Test II, when directly compared against gold-reference samples, our generated reasoning was still preferred in 39.7% of cases (95% CIs: [30.2%, 49.2%]). This finding suggests that providing explicit reasoning alongside emotion predictions can sometimes outweigh strict alignment with majority-vote labels. In other words, human raters may value the plausibility and interpretability of the reasoning process itself, highlighting the importance of considering minority labels as correct under our proposed explainable modeling framework.

### v2

@@ 1 Introduction
These limitations are most acute at evaluation time. Majority-only metrics penalize predictions that agree with well-supported but non-majority annotations, thereby underestimating performance on inherently ambiguous utterances and masking where models align with valid minority judgments [23]. To address this, SER research has increasingly embraced soft-label formulations, where the full annotator distribution or uncertainty-aware targets are preserved in training and evaluation [28, 8, 13]. Such methods not only yield higher accuracy on ambiguous cases but also align more faithfully with human variability. Beyond scores, a model requires understanding how systems map linguistic content and acoustic cues (prosody, pitch, energy, timing) to emotions so we can target weaknesses (e.g., lexical shortcuts vs. prosodic evidence). While recent SpeechLLM studies have begun entering the SER domain [21, 27, 5, 1, 18], and some report headline accuracies in technical reports [12], the majority still treat these models primarily as classifiers. Though Chen et al. [7] improve explainable SER via fine-grained speech emotion descriptors, descriptors are predefined from the pretrained model. This paradigm under-utilizes their generative capabilities, leaving untapped the potential to combine reasoning, explanation, and label uncertainty into a more holistic framework for emotion recognition.
In this work, we introduce an explainable SpeechLM framework that outputs a natural-language rationale alongside the emotion label. A reasoning-capable teacher LLM synthesizes concise explanations grounded in both lexical content and acoustic cues simultaneously. These rationales serve as intermediate supervision signals for a SpeechLM backbone. This teacher–student rationale supervision is shown effective in NLP that distilling teacher-generated reasoning traces can transfer rationalization capabilities to student models [16]. Given an utterance, the model first emits an ASR transcript to preserve transcription skill, then an emotion decision followed by a brief evidence-based rationale, all within a single coherent response. This structure encourages the model to provide both the emotion prediction and its rationale, improving interpretability while leaving majority-label training unchanged. To our knowledge, this is the first SER study to pair qualitative/quantitative rationale analysis with annotator-aware evaluation on a large in-the-wild corpus.
@@ 2.1 Reasoning Generation
Recent advances in reasoning-capable LLMs have made strong open models widely available. To bridge the gap between speech inputs and categorical emotion outputs, we leverage these models to generate human-readable rationales for each input–output pair in speech emotion recognition (SER). Note that rationales are conditioned on the gold label and thus reflect a label-informed explanation. The prompt template used for rationale synthesis is provided in the Prompt Design box, where the \$CONTENT and \$EMOTION mean the transcript and ground-truth emotion labels of the training utterances respectively.
@@
Concretely, we adopt Qwen2.5-Omni as the reasoning agent due to its strong multimodal understanding and reasoning capabilities. For each utterance, we provide the audio signal (encoded into embeddings by audio encoder) concatenated with prompt token embeddings consisting of its transcript and the reference emotion label to generate rationales for training data. Note that unlike prior work, we do not discretize acoustic cues into fixed levels. This strategy enables diverse natural-language descriptions and reflects real-world settings where the target distribution and class granularity are unknown. All rationales are generated exclusively on the training split, ensuring no label information leaks into validation or test evaluation.
@@ 2.2 Supervised Fine-Tuning (SFT)
Table 1: Baseline model comparison with Macro-F1 score on the MSP-PODCAST testing set ; denotes p-value $\leq$ 0.05.
@@
11 Q2A-lora-SFT w/ reasoning $\textbf{36.14\%}^{\star}$ $\textbf{66.78\%}^{\star}$ $\textbf{40.15\%}^{\star}$ $\textbf{59.32\%}^{\star}$ $\textbf{30.59\%}^{\star}$ $\textbf{62.55\%}^{\star}$ $\textbf{33.64\%}^{\star}$ $\textbf{54.13\%}^{\star}$
@@ 3.2 Baseline Comparisons
In contrast, our proposed SFT-based model (Model 11) substantially outperforms all zero-shot baselines (including CoT-based models), reaching 36.1% Macro-F1 on MV-labels (open-form) and 30.6% in the closed-form setting. Importantly, these gains are not merely a byproduct of fine-tuning: traditional SFT models trained directly for SER (Models 1–3) achieve only 23–30% on closed-form metrics, highlighting that simply applying supervised training does not guarantee strong performance. Instead, the improvement of Model 11 stems from incorporating rationale supervision during fine-tuning, which enhances both predictive accuracy and interpretability.
3.3 Ablation on Rationales
@@ 3.4 Subjective Test
To assess the reasoning plausibility, we selected $32$ utterances where the model’s predicted label disagreed with the reference. For each utterance, we formed two candidates including Model-predicted and Gold-referenced (rationale generated by our prompting procedure using the reference label as guidance). Raters marked the preference one, and we report the Gold-vs-Predicted preference rate.
Figure 1 summarizes the results. In Test I, the overall average acceptance rate reached 69.1% (95% confidence intervals(CIs): [58.9%, 79.2%]), with 72.5% for the match condition and 65.6% for the non-match condition. These results indicate that the reasoning generated by our model is perceived as plausible by human raters, even when the predicted emotion does not align with the ground truth.
In Test II, our rationales were preferred over gold-referenced ones in 39.7% of cases (95% CIs: [30.2%, 49.2%]). This finding suggests that providing explicit reasoning alongside emotion predictions can sometimes outweigh strict alignment with majority-vote labels. In other words, human raters may value the plausibility and interpretability of the reasoning process itself, highlighting the importance of considering minority labels as correct under our proposed explainable modeling framework.